OpenStudy (smored):

The roots of 2x^2-x+k=0 are sinθ and cosθ. Find k.

OpenStudy (owlfred):

\[\sin \theta + \cos \theta = \frac{1}{2}\tag{*}\]You need to find \(k\), and you know that \(\frac{k}{2}\) is the product of roots, meaning that \(\frac{k}{2} = \sin\theta \cdot \cos \theta\Rightarrow k = 2\sin \theta \cos \theta\). So you need to find \(2\sin\theta \cos \theta\). You can do so by squaring \((*)\).

OpenStudy (smored):

\[\sinθ+\cosθ=\frac{1}{2}\]??? really? im pretty sure its\[\sin^2θ+\cos^2θ=1\]

OpenStudy (perl):

this is a good resource for the sum and products of the roots of a quadratic polynomial http://www.mathsisfun.com/algebra/polynomials-sums-products-roots.html

OpenStudy (perl):

for a quadratic polynomial ax^2 + bx + c let r1, r2 be the roots . Then we have r1 + r2 = -b/a r1 * r2 = c/a

OpenStudy (smored):

at some point, do i need to find theta?

OpenStudy (owlfred):

None.

OpenStudy (owlfred):

If you square both sides in \(\sin \theta + \cos \theta = \frac{1}{2}\), which I got from Vieta's Formulas, what would you get?

OpenStudy (smored):

ahh

OpenStudy (smored):

thanks.

OpenStudy (owlfred):

Good to see that. :) What do you get?

OpenStudy (smored):

-.75